Individuals filing federal income tax returns prior to March 31 received an average refund of $1,077. Consider the population of last-minute filers who mail their tax return during the last five days of the income tax period (typically April 10 to April 15 a. A researcher suggests that a reason individuals wait until the last five days is that on average these individuals receive lower refunds than do early filers Develop appropriate hypotheses such that rejection of Ho will support the researchers contention. Ho: μ is equal to 1,077 : μ is less than 1,077 b. For a sample of 400 individuals who filed a tax return between April 10 and 15, the sample mean refund was $920. Based on prior experience a population standard deviation of ơ $1,800 may be assumed, what is the p-value (to 4 decimals)? c. Using α = 0.05, can you conclude that the population mean refund for last minute filers is less than the population mean refund for early filers? -Select your answer-▼ Answer the next three questions using the critical value approach d. Using α = 0.05, what is the critical value for the test statistic? Enter negative value as negative number State the rejection rule: Reject α = 0.05 ífz isl -select your answer. ▼| the critical value Using α = 0.05, can you conclude that the population mean refund for last minute filers is less than the population mean refund for early filers? - Select your answer -
Answer #1

b)

standard errror of mean = σx=σ/√n= 90.0000
test statistic           z             =    (xbar-μ)/σx= -1.744

hence p value =0.0409 ( please try 0.0405 if this comes wrong due to rounding error)

c)

No

d)

critical value of test statistic =-1.645

reject if z is less then the critical value

using alpha=0.05 , can you conclude,,,,,,,,,,,,? -No